OpenStudy (anonymous):

What's the rate of change and the y-intercept of y=7-6x

OpenStudy (anonymous):

The y-intercept the equation is 7.

OpenStudy (anonymous):

Isn't it y=m-bx

OpenStudy (astrophysics):

rate of change is the slope

OpenStudy (astrophysics):

and y - intercept is where the line crosses y - axis y=mx+b

OpenStudy (astrophysics):

m = slope b = y - int

OpenStudy (anonymous):

This is because the equation y=7-6x can also be expressed as y=-6x+7, which is the form of y=mx+c, where m is the gradient / slope / rate of change and c is the y-intercept

OpenStudy (mathmale):

You have y=7-6x and can and probably should re-write that as y=-6x+7. Comparing this to

OpenStudy (mathmale):

y=mx+b, slope m=-6 and y-intercept b=7.

OpenStudy (mathmale):

y=m-bx? unfortnately, no.

OpenStudy (mathmale):

Your y=7-6x has the form y=b + mx, with b=7 and m=-6.